# The Road to Trust: Building Enclaves within Confidential VMs

### v2

@@ 3.2. Enclave life cycle management
NestedSGX achieves compatibility with Intel SGX by transitioning to the security monitor on these leaf functions. Since VMPL switches can only be executed in the privileged mode (Sec. 2.2), NestedSGX provides the NestedSGX-driver within the guest OS, which accepts the request from the App and facilitates the VMPL switch. During this process, the NestedSGX-driver receives the type of SGX leaf functions, while the parameters are saved on the App’s stack. The NestedSGX-driver then switches the execution to VMPL0 and transfers control to the security monitor, where the security monitor emulates the instructions strictly adhering to the SGX specifications.
@@
Figure 3. Handling synchronous and asynchronous enclave entry and exit in NestedSGX.
@@
Synchronous enclave entry and exit. Once an enclave is initialized, the App can enter the enclave using the EENTER leaf function and jump to the enclave’s code for execution. As shown in Figure  3, during the emulation of EENTER, the security monitor switches to the enclave’s page table and enters the enclave entry point specified by the TCS using the sysret instruction. The enclave can then return to the App using the EEXIT leaf function, which is also emulated by the security monitor. Specifically, the enclave returns to the security monitor using the syscall instruction. Subsequently, the security monitor switches the execution to the NestedSGX-driver running at VMPL1. The NestedSGX-driver restores the App’s page table and returns control to the App using the sysret instruction.
@@
Asynchronous enclave entry and exit. During enclave execution, faults and exceptions may occur. In accordance with the SGX model, these faults and exceptions are supposed to be handled by the guest OS first. In NestedSGX, when a fault or exception occurs within the enclave, it traps to the fault and exception handler within the security monitor (Figure 3). The security monitor then emulates the Asynchronous Exit (AEX) event by saving the context of the enclave in the state saving area (SSA). Subsequently, the security monitor switches the vCPU to the NestedSGX-driver, which operates at VMPL1. The NestedSGX-driver fills the instruction pointer (RIP) to the trampoline area (AEP) and invokes the handler of the guest OS. After the OS handles the fault or exception, it uses the iret instruction to return to the AEP. The AEP, in turn, resumes the execution of the enclave using the ERESUME leaf function. The emulation of the ERESUME leaf function is similar to that of EENTER, except that the enclave’s execution context is restored from the SSA. As a result, NestedSGX is capable of handling exceptions that occur during enclave execution. For example, NestedSGX supports the emulation of the CPUID instruction, enabling the execution of unmodified applications. [^2] In comparison, vSGX (Zhao et al., 2022) requires modifying the application code in order to bypass the CPUID check.
@@
It’s worth noting that if the untrusted NestedSGX-driver chooses not to fill RIP with AEP, this decision does not introduce any new security issues. This is because the control flow outside the enclave is already vulnerable to manipulation within the SGX model and is not inherently trusted. In NestedSGX, faults and exceptions are firstly trapped to the security monitor, enabling easy detection of suspicious behaviors such as abnormal interrupts caused by page fault-based attacks. Conceptually, NestedSGX can also support the recent AEX-notify feature (Constable et al., 2023), although the implementation is planned for future work.
3.3. Memory isolation
@@
Figure 4. The guest physical address space is divided into 3 parts: the secure memory used by the security monitor, the EPC memory, and the normal memory.
@@
The entire guest physical address space of the CVM is divided into 3 parts (Figure 4): the secure memory used by the security monitor, the secure memory used by the enclaves (i.e., EPC), and the normal memory used by the guest OS and untrusted applications. The access to the secure memory is restricted solely to VMPL0, whereas the normal memory can be accessed by both VMPL0 and VMPL1. Notably, the SGX control structures are stored in the memory of the security monitor, while the enclave’s code and data pages are stored in the memory reserved for the EPC.
@@
Isolation of enclaves. It is also important to prevent an enclave from tampering with the security monitor or the guest OS. This ensures protection against potentially malicious enclave (i.e., enclave malware) or enclaves that may have vulnerabilities that could be exploited. One approach to address this concern is to place the enclave at a distinct VMPL level, separate from VMPL0 or VMPL1, and enforce the appropriate VMPL permissions to contain the enclave code. However, the transition is emulated by the security monitor, necessitating at least two VMPL switches. As a result, transitioning between the application and the enclave becomes costly.
@@
3.4. Measurement, attestation and sealing
@@
Figure 5. The NestedSGX attestation report consists of two parts: the SEV-SNP attestation report signed by VCEK, and the enclave report signed by NestedSGX AIK. The NestedSGX AIK is generated within the security monitor and subsequently bound to the SEV-SNP hardware by placing its public key as part of the SEV-SNP attestation report.
Chain of trust. When the CVM is first initialized, the security monitor randomly generates a key pair, which serves as the attestation identity key (AIK) of NestedSGX. The AIK is only required to be generated once and can be reused across different CVM boot cycles. To establish the binding between NestedSGX’s AIK and the CVM, we include the digest of the public part of the attestation key in the user-data field of the attestation report request (Figure 5), while the private key never leaves the memory of the security monitor.
@@
4. Implementations
Table 2. NestedSGX software stack: the security monitor extends the SVSM framework and is mostly written in Rust and assembly. The NestedSGX-driver and Intel SGX SDK are mostly written in C/C++ and assembly.
@@ 4. Implementations
Our implementation of NestedSGX is based on the software stack provided by AMD, which is open-sourced on GitHub (AMD, 2023a). The code consists of several components, including QEMU, Open Virtual Machine Firmware (OVMF), and Linux kernels for both host and guest environments. To implement NestedSGX, we extend the existing open-source Linux Secure VM Service Module (SVSM) framework (SVSM, 2022; AMD, 2023c), adding around 5,500 lines of code to the SVSM implementation. For VMPL switches, NestedSGX employs the standard Model Specific Register (MSR) protocol outlined in the Guest-Host Communication Block (GHCB) specification (AMD, 2023b). The MSR protocol notifies the KVM to switch to the targeted VMPL, without any modification of the KVM.
Security monitor. We based our implementation on the main branch of SVSM without current privilege level (CPL) support, and added our own support for running user-space enclaves. The security monitor emulates SGX leaf functions based on requests from the NestedSGX-driver. To accommodate these requests, a new type of SVSM protocol called SGX protocol was introduced within the SVSM framework (SVSM, 2022). The security monitor effectively enters a request loop within the SVSM framework, allowing it to continue execution upon the arrival of subsequent requests. The security monitor handles a memory pool acting as the EPC, and faithfully emulates SGX data structures and operations in strict adherence to the SGX specification. We registered customized fault and exception handlers for the enclave, emulating the AEX event. Furthermore, we configured the system call handler by setting the LSTAR MSR to handle the EEXIT leaf function.
To enter the enclave, the security monitor performs a series of actions. It switches to the enclave’s page table and stack, and then executes the sysret instruction with RCX pointing to the entry address, effectively emulating the behavior of EENTER and ERESUME leaf functions. The EEXIT leaf function, on the other hand, is replaced with the syscall instruction. In this case, the system call handler within the security monitor manages the VMPL switches. In order to handle AEX events, the security monitor directly accesses the enclave’s State Saving Area (SSA) by disabling Supervisor Mode Access Prevention (SMAP). This allows the security monitor to handle AEX events efficiently. Since the security monitor and the enclaves have separate page tables, we put the privilege switching code in a trampoline page that is mapped in both page tables, which is similar to the implementation of kernel page table isolation (KPTI) (Gruss et al., 2017).
@@
With the aforementioned design, the majority of SGX programs can run on NestedSGX without requiring modifications to the source code. All the sample code projects provided by the official Intel SGX SDK, except SampleAEXNotify, ran without any issues on NestedSGX. Furthermore, to simplify the porting of existing applications, we have adapted the Rust SGX SDK (Wang et al., 2019) to the NestedSGX platform. Furthermore, we made slight modifications to the Occlum library OS (Shen et al., 2020) to handle the marshalling buffer and enable its execution on NestedSGX. Building upon Occlum, we implemented the HotCalls optimization (Weisse et al., 2017), allowing certain OCALLs to be processed asynchronously within the enclave without exiting it. [^3] We did not cover all OCALLs but focused on optimizing 15 frequently used ones, such as occlum_ocall_sendmsg, occlum_ocall_recvmsg, occlum_ocall_clock_gettime, occlum_ocall_posix_memalign, occlum_ocall_free, u_sgxprotectedfs_fread_node and u_setsockopt_ocall etc.
Intra-enclave isolation. Since Intel SGX only supports the monolithic model, attempts have been made to support intra-enclave isolation, such as Nested Enclave (Park et al., 2020) and LIGHTENCLAVE (Gu et al., 2022). However, these solutions necessitate hardware modifications that go beyond Intel SGX’s capabilities. We incorporated support for intra-enclave isolation, leveraging the hardware capabilities of memory protection keys for userspace (PKRU). We offer four PKRU APIs as syscalls for enclaves: pkey_set, pkey_alloc, pkey_free, and pkey_mprotect. Additionally, we provide two syscalls that allow enclaves to enable or disable PKRU mechanisms.
Initially, the enclaves invoke the pkru_enable function, informing the security monitor to establish the key manager. Subsequently, the enclaves can use the pkey_alloc and pkey_set functions to obtain a key from the manager and make use of it through pkey_mprotect. On the pkey_mprotect function, the security monitor checks with the key manager to verify if the key has been assigned to the enclaves. If valid, the monitor updates the PKRU-bits of corresponding PTEs by traversing the enclaves’ page tables. The key manager faithfully records and updates key information, including key attributes and the protected virtual address of enclaves. When the enclaves invoke the pkru_free function, the security monitor clears the PKRU-bits of the relevant PTEs. Additionally, the enclaves can call the pkru_disable function to instruct the security monitor to release the key manager.
5. Security Analysis
We conduct the security analysis by enumerating the attack vectors under the threat model outlined in Section 2.3, and describing how NestedSGX defends against them.
@@ 5. Security Analysis
Untrusted enclave. Although the enclave operates at the highest VMPL, it is not allowed to access the gPA belonging to lower VMPLs, such as the lower VMPL’s VMSA. This is achieved by page table based isolation. Only the enclave’s pages are mapped in its page table which is being managed by the security monitor. The hardware also prevents the enclave from overwriting the VMPL permissions of guest pages, since the instruction to do so (RMPADJUST, Sec. 2.2) is privileged and can only be executed by the security monitor.
@@
Chain of trust analysis. The NestedSGX report is comprised of the SEV-SNP attestation report and the enclave report. Any attempt to tamper with the security monitor will result in a change to the SEV-SNP attestation report, thereby ensuring that the security monitor is reliable and performs as expected. Specifically, the security monitor generates NestedSGX’s AIK and maintains its private key inaccessible to other components. It associates NestedSGX’s AIK with the platform by including the digest of the AIK’s public key in the SEV-SNP attestation report.
While the enclave operates at VMPL0, it is prohibited from directly obtaining the SEV-SNP attestation report. Otherwise, the enclave may incorporate a counterfeit AIK into the report, which could compromise the integrity of the trust chain, particularly if the enclave deliberately exposes the private key of the AIK. This protection is enforced by the fact that SNP_REPORT_REQ can only be initiated from the kernel mode, which is the security monitor. More specifically, SNP_REPORT_REQ necessitates shared memory with the host VMM for passing request and response messages, but we intentionally disable shared memory between the enclave and the host VMM. The hardware also prevents the guest OS (operating at VMPL1) from obtaining an SEV-SNP attestation report for VMPL0. Specifically, the desired VMPL is provided by the guest in the SNP_REPORT_REQ message, and the guest can only generate attestation reports for VMPLs that are greater than or equal to the current VMPL (AMD, 2023d, Sec. 7.3).
@@
6. Evaluations
6.1. Experimental setup
@@
The baseline was conducted on the same AMD server running in the SGX SDK simulation mode (sim mode), which does not provide any security guarantees. Specifically, we used the same code base under the same optimization level and compiling options. Notably, as a result of the HotCalls optimization, applications running on Occlum experienced improved performance without the need for any modifications to their source code.
6.2. Micro-benchmarks
[b]
Table 3. Latencies of emulated SGX leaf instructions. The emulation of ENCLS leaf instructions require VMPL switches, which dominates the cost, while most ENCLU leaf instructions are emulated entirely at VMPL0. The slower performance of ENCLU leaf instructions on NestedSGX is attributed to the use of slower Rust implementations in the cryptographic crates.
@@ 6.2. Micro-benchmarks
  • $\dagger$
    The vSGX results are taken from the published paper (Zhao et al., 2022).
First, we measured the latencies of emulated SGX leaf instructions and context switches (i.e., ECALLs and OCALLs). Following that, we conducted 4 representative micro-benchmarks: Linux/Unix nbench for memory- and computation-intensive tasks, Lmbench for memory-intensive tasks, WolfSSL for computation-intensive tasks, and FIO for I/O-intensive tasks.
SGX leaf instructions. We measured the emulated SGX leaf instructions in NestedSGX, averaging the results over 10,000 runs. As shown in Table 3, all ENCLS leaf instructions on NestedSGX are significantly faster than vSGX. This is attributed to the requirement of VMPL switches for emulating ENCLS leaf instructions on NestedSGX, whereas vSGX necessitates more costly cross-VM communications. Both EGETKEY and EREPORT on NestedSGX are slower than those on vSGX, which we confirmed is due to the use of the Rust rust-crypto crate that is slower than the implementation used by vSGX. Notably, EGETKEY and EREPORT are not heavily invoked in practice.
[b]
Table 4. Latencies of context switches. The cost of NestedSGX includes the following parts: a VMPL switch round-trip which costs about 19,400 cycles on our platform, the SGX SDK routine which costs 4000–5500 cycles, and the rest is the cost of our unoptimized implementation of the security monitor in emulating SGX instructions (including an EENTER and an EEXIT, which cost about 9,000 cycles). The Intel platform for comparison was equipped with an Intel Xeon E3-1270 v6 CPU and 64 GB RAM.
@@
ECALL 14,437 cycles (3.8 us) $\approx 1,500$ us 33,584 cycles (12 us)
OCALL 12,586 cycles (3.3 us) - 32,014 cycles (11 us)
  • $\ddagger$
    The vSGX results are taken from the published paper (Zhao et al., 2022).
Context switches. We measured the latency of ECALLs and OCALLs in NestedSGX. The test runs empty edge calls with no explicit parameters 100,000 times and takes the average value. As shown in Table 4, the latency of context switches in NestedSGX is 1.33×1.$3$1.54×1.54$\times$1.54 × higher than Intel SGX, but is still two orders of magnitude faster than vSGX. Notably, existing works on Intel SGX, including EleOS (Orenbach et al., 2017), HotCalls (Weisse et al., 2017) and Switchless calls (Tian et al., 2018) can be further applied to NestedSGX to reduce the cost brought by context switches.
@@
Figure 6. Micro-benchmarks.
Linux/Unix nbench. Linux/Unix nbench is a benchmark suite that focuses on evaluating the performance of a computer system’s CPU, FPU, and memory system (Mayer, 2017). We utilized an adapted version of nbench, namely SGX-NBench (sgx, 2017) to evaluate the computation performance of NestedSGX. Figure 6(a) demonstrates that NestedSGX introduces an overhead of approximately 1.3% on average over the baseline, which we believe mostly attributes to the context switches introduced by SGX-NBench during the evaluation.
Lmbench. Lmbench is a benchmark suite commonly employed to assess various aspects of system behavior. In our evaluation, we utilized the SGX version of Lmbench, known as SGX-bench (Hasan et al., 2020), specifically employing the bw_mem test to evaluate the memory bandwidth. We conducted 10,000 repetitions of the test and recorded the average values. The memory block size for every access was set to 2 MB. The results, presented in Figure 6(c), indicate that the memory bandwidth achieved is 98.7% of the baseline.
WolfSSL (wol, 2024a). We performed an evaluation of WolfSSL with Intel SGX (wol, 2024b). This benchmark primarily focuses on computation-intensive tasks, such as encryption, decryption, digests, and signature verification. As illustrated in Figure 6(d), NestedSGX introduces an average overhead of about 0.78% compared to the baseline.
Flexible I/O Tester (FIO) (git, 2024) with Occlum. We conducted an evaluation of the I/O performance of NestedSGX by utilizing FIO (v3.28) with Occlum in the default configuration. Our focus was on benchmarking the bandwidth of randomized and sequential read and write operations to the disk, utilizing a single thread. We chose the direct I/O mode for sequential operations. Each test involved accessing a 256 MB file with a block size of 256 KB per access. We repeated the test 10 times, with each iteration lasting 100 seconds, and calculated the average bandwidth from all the runs. As depicted in Figure 6(c), the I/O read performance of NestedSGX is approximately 65.3% of the baseline, while the I/O write performance is similar to the baseline, possibly due to Occlum’s optimizations to reduce the context switches in write operations.
6.3. Real-world evaluations

### v3

@@ III-B Enclave life cycle management
NestedSGX achieves compatibility with SGX by transitioning to the security monitor on these leaf functions. Since VMPL switches can only be executed in the privileged mode, NestedSGX provides the NestedSGX-driver within the guest OS, which accepts the request from the App and facilitates the VMPL switch. During this process, the NestedSGX-driver receives the type of SGX leaf functions, while the parameters are saved on the App’s stack. The NestedSGX-driver then switches the execution to VMPL0 and transfers control to the security monitor, where the security monitor emulates the instructions strictly adhering to the SGX specifications.
@@
Figure 3: The management of enclave life cycles, e.g., handling synchronous and asynchronous enclave entry and exit events.
@@
Synchronous enclave entry and exit. Once an enclave is initialized, the App can enter the enclave using the EENTER leaf function and jump to the enclave’s code for execution. As shown in Fig. 3, during the emulation of EENTER, the security monitor switches to the enclave’s page table and enters the enclave entry point specified by the TCS using the sysret instruction. The enclave can then return to the App using the EEXIT leaf function, which is also emulated by the security monitor. Specifically, the enclave returns to the security monitor using the syscall instruction. Subsequently, the security monitor switches the execution to the NestedSGX-driver running at VMPL1. The NestedSGX-driver restores the App’s page table and returns control to the App using the sysret instruction.
@@
Asynchronous enclave entry and exit. During enclave execution, faults and exceptions may occur. According to the SGX model, these faults and exceptions are supposed to be handled by the guest OS first. In NestedSGX, when a fault or exception occurs within the enclave, it traps to the fault and exception handler within the security monitor (Fig. 3). The security monitor then emulates the Asynchronous Exit (AEX) event by saving the context of the enclave in the state saving area (SSA). The security monitor switches the vCPU to the NestedSGX-driver, which operates at VMPL1. The NestedSGX-driver fills the instruction pointer (RIP) to the trampoline area (AEP) and invokes the handler of the guest OS. After the OS handles the fault or exception, it uses the iret instruction to return to the AEP. The AEP, in turn, resumes the enclave execution with the ERESUME leaf function. The emulation of the ERESUME leaf function is similar to that of EENTER, except that the enclave’s execution context is restored from the SSA. As a result, NestedSGX is able to handle exceptions that occur during enclave execution. For example, NestedSGX supports the emulation of the CPUID instruction, enabling the execution of unmodified applications.  [^2] In comparison, vSGX [62] requires modifying the application code in order to bypass the CPUID check.
@@
If the untrusted NestedSGX-driver chooses not to fill RIP with AEP, this decision does not introduce any new security issues. This is because the control flow outside the enclave is already vulnerable to manipulation within the SGX model and is not inherently trusted. In NestedSGX, faults and exceptions are firstly trapped to the security monitor, enabling easy detection of suspicious behaviors such as abnormal interrupts caused by page fault-based attacks. Conceptually, NestedSGX can also support the recent AEX-notify feature [23], although the implementation is planned for future work.
III-C Memory isolation
@@
Figure 4: The guest physical address space is divided into 3 parts: the secure memory used by the security monitor, the EPC memory, and the normal memory.
@@
The entire guest physical address space of the CVM is divided into 3 parts (Fig. 4): the secure memory used by the security monitor, the secure memory used by the enclaves (i.e., EPC), and the normal memory used by the guest OS and untrusted applications. The access to the secure memory is restricted solely to VMPL0, whereas the normal memory can be accessed by both VMPL0 and VMPL1. Notably, the SGX control structures are stored in the memory of the security monitor, while the enclave’s code and data pages are stored in the memory reserved for the EPC.
@@
Isolation of enclaves. It is also important to prevent an enclave from tampering with the security monitor or the guest OS. This ensures protection against potentially malicious enclave (i.e., enclave malware) or enclaves that may have vulnerabilities that could be exploited. One approach to address this concern is to place the enclave at a distinct VMPL level, separate from VMPL0 or VMPL1, and enforce the appropriate VMPL permissions to contain the enclave code. However, the transition is mediated by the security monitor, necessitating at least two VMPL switches. As a result, transitioning between the application and the enclave becomes costly.
@@
III-D Measurement, attestation and sealing
@@
Figure 5: The NestedSGX attestation report consists of two parts: the SEV-SNP attestation report signed by VCEK, and the enclave report signed by NestedSGX AIK. The NestedSGX AIK is generated within the security monitor and subsequently bound to the SEV-SNP hardware by placing its public key as part of the SEV-SNP attestation report.
Chain of trust. When the CVM is first initialized, the security monitor randomly generates a key pair, which serves as the attestation identity key (AIK) of NestedSGX. The AIK is only required to be generated once and can be reused across different CVM boot cycles. To establish the binding between NestedSGX’s AIK and the CVM, we include the digest of the public part of the attestation key in the user-data field of the attestation report request (Fig. 5), while the private key never leaves the memory of the security monitor.
@@
IV Implementations
Table II: NestedSGX software stack: the security monitor extends the SVSM framework and is mostly written in Rust and assembly. The NestedSGX-driver and Intel SGX SDK are mostly written in C/C++ and assembly.
@@ IV Implementations
Our implementation of NestedSGX is based on the software stack provided by AMD, which is open-sourced on GitHub [14]. The code consists of several components, including QEMU, Open Virtual Machine Firmware (OVMF), and Linux kernels for both host and guest environments. To implement NestedSGX, we extend the existing open-source Linux Secure VM Service Module (SVSM) framework [50, 16], adding around 5,500 lines of code to the SVSM implementation. For VMPL switches, NestedSGX employs the standard Model Specific Register (MSR) protocol outlined in the Guest-Host Communication Block (GHCB) specification [15]. The MSR protocol notifies the KVM to switch to the targeted VMPL, without any modification of the KVM.
Security monitor. We based our implementation on the main branch of SVSM without CPL3 support, and added our own support for running user-space enclaves. The security monitor emulates SGX leaf functions based on requests from the NestedSGX-driver. To accommodate these requests, a new type of SVSM protocol called SGX protocol was introduced within the SVSM framework [50]. The security monitor effectively enters a request loop within the SVSM framework, allowing it to continue execution upon the arrival of subsequent requests. The security monitor handles a memory pool acting as the EPC, and faithfully emulates SGX data structures and operations in strict adherence to the SGX specification. We registered customized fault and exception handlers for the enclave, emulating the AEX event. Furthermore, we configured the system call handler by setting the LSTAR MSR to handle the EEXIT leaf function.
To enter the enclave, the security monitor performs a series of actions. It switches to the enclave’s page table and stack, and then executes the sysret instruction with RCX pointing to the entry address, effectively emulating the behavior of EENTER and ERESUME leaf functions. The EEXIT leaf function, on the other hand, is replaced with the syscall instruction. In this case, the system call handler within the security monitor manages the VMPL switches. In order to handle AEX events, the security monitor directly accesses the enclave’s State Saving Area (SSA) by disabling Supervisor Mode Access Prevention (SMAP). This allows the security monitor to handle AEX events efficiently. Since the security monitor and the enclaves have separate page tables, we put the privilege switching code in a trampoline page that is mapped in both page tables, which is similar to the implementation of kernel page table isolation (KPTI) [26].
@@
With the aforementioned design, the majority of SGX programs can run on NestedSGX without requiring source code modifications. All the sample code projects provided by the official Intel SGX SDK, except SampleAEXNotify, ran without any issues on NestedSGX. Furthermore, we have adapted the Rust SGX SDK [56] and the Occlum library OS [48] to the NestedSGX platform. Building upon Occlum, we implemented the HotCalls optimization [58], allowing certain OCALLs to be processed asynchronously within the enclave without exiting it. [^3] We did not cover all OCALLs but focused on optimizing 15 frequently used ones, such as occlum_ocall_sendmsg, occlum_ocall_recvmsg, occlum_ocall_clock_gettime, occlum_ocall_posix_memalign, occlum_ocall_free, u_sgxprotectedfs_fread_node and u_setsockopt_ocall etc.
Intra-enclave isolation. Since Intel SGX only supports the monolithic model, attempts have been made to support intra-enclave isolation, such as Nested Enclave [42] and LIGHTENCLAVE [27]. However, these solutions necessitate hardware modifications that go beyond Intel SGX’s capabilities. We incorporated support for intra-enclave isolation, leveraging the hardware capabilities of memory protection keys for userspace (PKRU). We offer four PKRU APIs as syscalls for enclaves: pkey_set, pkey_alloc, pkey_free, and pkey_mprotect. Additionally, we provide two syscalls that allow enclaves to enable or disable PKRU mechanisms.
Initially, the enclaves invoke the pkru_enable function, informing the security monitor to establish the key manager. Subsequently, the enclaves can use the pkey_alloc and pkey_set functions to obtain a key from the manager and make use of it through pkey_mprotect. On the pkey_mprotect function, the security monitor checks with the key manager to verify if the key has been assigned to the enclaves. If valid, the monitor updates the PKRU-bits of corresponding PTEs by traversing the enclaves’ page tables. The key manager faithfully records and updates key information, including key attributes and the protected virtual address of enclaves. When the enclaves invoke the pkru_free function, the security monitor clears the PKRU-bits of the relevant PTEs. Lastly, the enclaves can call the pkru_disable function to instruct the security monitor to release the key manager.
V Security Analysis
We conduct the security analysis by enumerating the attack vectors under the threat model outlined in Section II-C, and describing how NestedSGX defends against them.
@@ V Security Analysis
Untrusted enclave. Although the enclave operates at the highest VMPL, it is not allowed to access the gPA belonging to lower VMPLs, such as the lower VMPL’s VMSA. This is achieved by page table based isolation. Only the enclave’s pages are mapped in its page table which is being managed by the security monitor. The hardware also prevents the enclave from overwriting the VMPL permissions of guest pages, since the instruction to do so (RMPADJUST, Sec. II-B) is privileged and can only be executed by the security monitor.
@@
Chain of trust analysis. The NestedSGX report includes the SEV-SNP attestation report and enclave report. Any attempt to tamper with the security monitor will result in a change to the SEV-SNP attestation report, thereby ensuring the security monitor is reliable and performs as expected. Specifically, the security monitor generates NestedSGX’s AIK and maintains its private key inaccessible to other components. It associates NestedSGX’s AIK with the platform by including the digest of the AIK’s public key in the SEV-SNP attestation report.
While the enclave operates at VMPL0, it is prohibited from directly obtaining the SEV-SNP attestation report. Otherwise, the enclave may incorporate a counterfeit AIK into the report, which could compromise the integrity of the trust chain, particularly if the enclave deliberately exposes the private key of the AIK. This protection is enforced by the fact that SNP_REPORT_REQ can only be initiated from the kernel mode, which is the security monitor. More specifically, SNP_REPORT_REQ necessitates shared memory with the host VMM for passing request and response messages, but we intentionally disable shared memory between the enclave and the host VMM. The hardware also prevents the guest OS (operating at VMPL1) from obtaining an SEV-SNP attestation report for VMPL0. Specifically, the desired VMPL is provided by the guest in the SNP_REPORT_REQ message, and the guest can only generate attestation reports for VMPLs that are greater than or equal to the current VMPL [17, Sec. 7.3].
@@
VI Evaluations
VI-A Experimental setup
@@
The baseline was conducted on the same AMD server running in the SGX SDK simulation mode (sim mode), which does not provide any security guarantees. Specifically, we used the same code base under the same optimization level and compiling options. Notably, as a result of the HotCalls optimization, applications running on Occlum experienced improved performance without the need for any modifications to their source code. We were not able to reproduce the experiment reported in vSGX, and the vSGX results are taken directly from the published paper [62]. The Intel platform for comparison was equipped with an Intel Xeon Platinum 8369B CPU and 64 GB RAM.
VI-B Micro-benchmarks
Table III: Latencies of emulated SGX leaf instructions. The emulation of ENCLS leaf instructions require VMPL switches, which dominates the cost, while most ENCLU leaf instructions are emulated entirely at VMPL0. The slower performance of ENCLU leaf instructions on NestedSGX is attributed to the use of slower Rust implementations in the cryptographic crates.
@@ VI-B Micro-benchmarks
SGX leaf instructions. We measured the emulated SGX leaf instructions in NestedSGX, averaging the results over 10,000 runs. As shown in Table III, all ENCLS leaf instructions on NestedSGX are significantly faster than vSGX. This is attributed to the requirement of VMPL switches for emulating ENCLS leaf instructions on NestedSGX, whereas vSGX necessitates more costly cross-VM communications. Both EGETKEY and EREPORT on NestedSGX are slower than those on vSGX, which we confirmed is due to the use of the Rust rust-crypto crate that is slower than the implementation used by vSGX. Notably, EGETKEY and EREPORT are not heavily invoked in practice.
Table IV: Latencies of context switches. The cost of NestedSGX includes the following parts: a VMPL switch round-trip which costs about 19,400 cycles on our platform, the SGX SDK routine which costs 4000–5500 cycles, and the rest is the cost of our unoptimized implementation of the security monitor in emulating SGX instructions (including an EENTER and an EEXIT, which cost about 9,000 cycles).
@@
ECALL 10,988 cycles (4.1 us) $\approx 1,500$ us 33,584 cycles (12 us)
OCALL 9,337 cycles (3.5 us) - 32,014 cycles (11 us)
Context switches. We measured the latency of ECALLs and OCALLs in NestedSGX. The test runs empty edge calls with no explicit parameters 100,000 times and takes the average value. As shown in Table IV, the latency of context switches in NestedSGX is 1.9×1.9$\times$1.9 ×2.1×2.1$\times$2.1 × higher than Intel SGX, but is still two orders of magnitude faster than vSGX. Notably, existing works on Intel SGX, including EleOS [41], HotCalls [58] and Switchless calls [51] can be further applied to NestedSGX to reduce the cost brought by context switches.
@@
Figure 6: Micro-benchmarks.
Linux/Unix nbench. Linux/Unix nbench is a benchmark suite that focuses on evaluating the performance of a computer system’s CPU, FPU, and memory system [38]. We utilized an adapted version of nbench, namely SGX-NBench [1] to evaluate the computation performance of NestedSGX. Fig. 6(a) demonstrates that NestedSGX introduces an overhead of approximately 1.3% on average over the baseline, which we believe mostly attributes to the context switches introduced by SGX-NBench during the evaluation.
Lmbench. Lmbench is a benchmark suite commonly employed to assess various aspects of system behavior. In our evaluation, we utilized the SGX version of Lmbench, known as SGX-bench [28], specifically employing the bw_mem test to evaluate the memory bandwidth. We conducted 10,000 repetitions of the test and recorded the average values. The memory block size for every access was set to 2 MB. The results, presented in Fig. 6(c), indicate that the memory bandwidth achieved is 98.7% of the baseline.
WolfSSL [11]. We performed an evaluation of WolfSSL with Intel SGX [12]. This benchmark primarily focuses on computation-intensive tasks, such as encryption, decryption, digests, and signature verification. As illustrated in Fig. 6(d), NestedSGX introduces an average overhead of about 0.78% compared to the baseline.
Flexible I/O Tester (FIO) [10] with Occlum. We conducted an evaluation of the I/O performance of NestedSGX by utilizing FIO (v3.28) with Occlum in the default configuration. Our focus was on benchmarking the bandwidth of randomized and sequential read and write operations to the disk, utilizing a single thread. We chose the direct I/O mode for sequential operations. Each test involved accessing a 256 MB file with a block size of 256 KB per access. We repeated the test 10 times, with each iteration lasting 100 seconds, and calculated the average bandwidth from all the runs. As depicted in Fig. 6(c), the I/O read performance of NestedSGX is approximately 65.3% of the baseline, while the I/O write performance is similar to the baseline, possibly due to Occlum’s optimizations to reduce the context switches in write operations.
VI-C Real-world evaluations